Suppose an economy has 10,000 people who are not working but looking and available for work and 90,000 people who are working. What is its unemployment rate?

Now suppose 4,000 of the people looking for work get discouraged and give up their searches. What happens to the unemployment rate? Would you interpret this as good news for the economy or bad news? Explain.


1
Expert's answer
2021-10-04T10:19:47-0400

"UnemploymentRate=\frac{Unemployed People}{LabourForce}\times100\%"

The unemployment rate shows the share of people in the labor force who do not currently have a job but are actively looking for it.

"UnemploymentRate=\frac{10,000}{90,000+10,000}\times100\%=10\%"

When workers become discouraged and stop looking for employment, they leave the labor force.  So, the unemployment rate will decrease:

"UnemploymentRate=\frac{10,000-4,000}{90,000+10,000-4,000}\times100\%=6.25\%"